OpenStudy (anonymous):

Evaluate 15/j when j =5. 3 10 20 75

OpenStudy (callisto):

15/j, when j=5 15/j = 15/5 =?

OpenStudy (anonymous):

75?

sam (.sam.):

15/5 means 15 divided by 5

OpenStudy (anonymous):

3

sam (.sam.):

yes

OpenStudy (anonymous):

tnxs so much

sam (.sam.):

np ;)